OpenStudy (jellybot23):

No one is in Physics, so I am posting this here, can anyone help? Compare the equation for a line through the origin, y = mx, to the equation Fc= mv^2/r, and explain what the slope of the graph of Fc vs. v2 represents. Remember that the m in the equation for a line represents the slope, and the m in the centripetal force equation represents the mass of the stopper. I don't understand what they are asking

OpenStudy (nuttyliaczar):

Well the slope for the force equation in terms of v^2 is going to be m/r, similar to how the slope for the y=mx+b is just m (the coefficient in front of the x). I think that's what it was asking about right?

OpenStudy (jellybot23):

Ohhhh, that makes sense! Thanks so much! I believe that is what they are looking for :)

OpenStudy (nuttyliaczar):

Yep no problem. Physics major so I would hope that I can help at least a little

OpenStudy (nuttyliaczar):

Although this was more math than anything else

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=mx }\) is a line throught the origin with positive slope m \(\color{#000000 }{ \displaystyle F_c=mv^2/r=ma }\) is the equation of the cetripetal force in terms of centripetal acceleration, where \(v^2/r=a\) (\(a\)-centripetal acceleration). they are both lines